Calculate the value of the translational partition function of O_2
2

at 1000 K for a volume of 1 m^3
3
. Submit your answer multiplied by 10^{-32}
−32
(i.e. if you calculate 4.2e33, input 42).
1
Expert's answer
2020-06-30T07:46:52-0400

Answer

Mass of O2 is 2 x 32 x 1.6606 x 10-27 kg

2πmkBT = 2 x 3.1415 x (2 x 32 x 1.6606 x 10-27kg) x 1.3807 x 10-23 J/K x 1000 K = 9.2196 x 10-45 J kg

Λ = h/(2πmkBT)1/2 = 6.6262 x 10-34 JS/(9.2196 x 10-45 J kg)1/2 = 6.90 x 10-12 m

The value of translational partition function = v/ Λ3 = 1m3/(6.90 x 10-12 m)3 = 30.4 x 1032

Multiplied by 10-32  the answer is 30.4 (Ans.)